OpenStudy (anonymous):

4 + 2/xy ALL OVER 2 simplify

OpenStudy (anonymous):

how do i begin?

OpenStudy (anonymous):



OpenStudy (anonymous):

Are you simplifying this expression?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[\frac{ 4 }{ 2 }+\frac{ 2 }{ 2xy } = 2+\frac{ 1 }{ xy }\]

OpenStudy (anonymous):

or \[\frac{ 2xy+1 }{ xy }\]

OpenStudy (anonymous):

how did you get that

OpenStudy (anonymous):

\[\frac{ 4+\frac{ 2 }{ xy } }{ 2 } = \frac{ 1 }{ 2 }(4+\frac{ 2 }{ xy })\] Distribute \[\frac{ 4 }{ 2 }+\frac{ 2 }{ 2xy }\]

OpenStudy (anonymous):

Make sense?

OpenStudy (anonymous):

yes thank you